OpenStudy (anonymous):

A square with sides of is inscribed in a circle. What is the area of one of the sectors formed by the radii to the vertices of the square? 1.5 2.25 4.5

OpenStudy (anonymous):

@Lilium

OpenStudy (anonymous):

Hence, (3sqrt(2))^2 + (3sqrt(2))^2 = d^2 Hence, d=6 (you can do the math following the equation above) so divide this by 2 in order to get the radius of the circle. therefore, r = 3 Next, you will notice that the coloured area we are looking for is only a quarter of the entire circle minus the area of a quarter of the entire square. So knowing that, first, you must the area of the circle using the equation: pi*r^2 = pi*(3)^2 = 9pi you need a quarter of that so divide it by 4 to get 9pi/4. Next, find the area of the square. length x width. So, 2*(3sqrt(2)) = 6sqrt(2) but similarly, you only need a quarter of that, so divide that by 4 to get 3sqrt(2)/2 Therefore, your final answer of the area of the coloured region will be the subtraction of the two areas: [(9pi/4)-(3sqrt(2)/2)] = (9pi-6sqrt(2))/4

OpenStudy (anonymous):

hope i helped :)

OpenStudy (anonymous):

4.5

OpenStudy (anonymous):

thanks